# Collusion-Resilience in Transaction Fee Mechanism Design

### v2

@@ 5.2 Formal Proofs
We now show that the confirmation probability under a single bid $a$ for a sufficiently large $a$ is non-zero assuming the miner honestly follows the inclusion rule. Suppose this is not true for the sake of a contradiction; that is, when the bid vector is just $a$ and the miner is honest, no one has positive confirmation probability, so the social welfare is zero. Since the mechanism is non-trivial, there must exist a scenario ${\mathbf{b}}=(b_{1},\dots,b_{t})$ such that some user $i$ has positive confirmation probability. Let $a>\sum_{j=1}^{t}b_{j}$. Then, imagine that there is only one user with true value $a$. The global coalition can replace the user’s primary bid with $b_{i}$, inject the fake bids ${\mathbf{b}}_{-i}$, and pretend that the bid vector is ${\mathbf{b}}$. This way, the user would get positive confirmation probability and since its payment is at most $\sum_{j=1}^{t}b_{j}$, the user’s expected utility (and thus the social welfare) is positive. This violates global SCP.
@@ 6.1 A Non-Truthful Mechanism with UIC + MIC + OCA-Proof
The rationale of the design is to signal to the mechanism when everyone is adopting the globally optimal strategy $\sigma$ (as opposed to the bidding rule used to establish UIC). When the mechanism detects that everyone is behaving according to $\sigma$, it adopts a different behavior to optimize social welfare. We use the range $[0,1)$ to encode the actual bid, and use the range $[1,\infty)$ for signalling. While the resulting mechanism is somewhat contrived and not necessarily meaningful from a practical point of view, it clarifies which notions of collusion-resilience most accurately capture the intended modeling goals and illustrates some technical challenges involved in the proof in Section 6.2. Consider the following TFM:
@@ 6.2 Impossibility of UIC + MIC + OCA-Proof for Truthful Mechanisms
The structure of the proof in this section is similar to the proof in Section 5. In fact, Lemmas 6.2, 6.3, 6.6, 6.7, 6.8 and 6.9 are the analogs of Lemmas 5.1, 5.2, 5.3, 5.4, 5.5 and 5.6, respectively, except that we need to work on the images of $\sigma$ in order to apply OCA-proofness. Before diving into the proof, we make a few remarks.
@@
  • As we have seen a feasibility for non-truthful mechanisms in Section 6.1, we must require the bidding rule to be truth-telling to reach an impossibility. Indeed, the proofs of Lemmas 6.4, 6.7, 6.8 and 6.9 rely on Myerson’s lemma, and Myerson’s lemma requires the mechanism to be truthful. Notice that the mechanism in Section 6.1 does not satisfy the requirements of Myerson’s lemma. If we apply the revelation principle to make the mechanism become truthful so that Myerson’s lemma could apply, it breaks OCA-proofness (see Remark 1).
@@
Imagine that $k+1$ players each have true values $b_{1},\ldots,b_{k+1}$, and they all bid according to the globally optimal strategy $\sigma$. It suffices to show the following: suppose $\sigma(b_{k+1})$ has a non-zero probability $p$ of being included in the block, then conditioned on being included, its probability of confirmation must be 00. Suppose this is not true, and conditioned on being included, $\sigma(b_{k+1})$ has probability $p^{\prime}>0$ of being confirmed. Then, the global coalition can adopt the following OCA: run the honest inclusion rule in its head, including flipping any coins required by the inclusion rule, and let ${\bf B}$ denote the resulting block. If $\sigma(b_{k+1})$ is not included in ${\bf B}$, simply output the block ${\bf B}$. Else if $\sigma(b_{k+1})$ is included in ${\bf B}$, then let $\sigma(b_{i})$ be some bid that is left out — such a bid must exist because the block size is only $k$. Now, have the $i$-th user bid $\sigma(b_{k+1})$ and make an actual block ${\bf B}$ where the bid $\sigma(b_{k+1})$ corresponds to user $i$’s bid. The expected increase in social welfare relative to everyone adopting $\sigma$ is $p\cdot p^{\prime}\cdot(b_{i}-b_{k+1})>0$, which violates OCA-proofness and the fact that $\sigma$ is the social-welfare-maximizing strategy. ∎
@@
A well-known fact is that any function that is monotone must have only countably many discontinuities within any interval $[x_{1},x_{2}]$ where $x_{2}>x_{1}$; further, all such discontinuities must be jump discontinuities. By Lemma 6.4, there exists a constant $c^{*}$ such that $\sigma$ is monotone for all inputs larger than $c^{*}$. Thus, for any $a>c^{*}$, within the interval $[a,a+1]$, $\sigma$ is monotone and there must exist a point $a^{*}$ such that $\sigma$ is continuous at $a^{*}$. ∎
@@
  • Moreover, if there are $k+1$ users with the true value $\mathbf{s}$ and $\sigma$ is continuous at $b$, the expected total user utility is 00 when all users bid truthfully.
@@
The base case is when $i=k$; that is, there are $k+1$ users with the true values $\mathbf{s}=(\underbrace{\sigma(b),\ldots,\sigma(b)}_{k},\sigma(b_{k+1}))$. Notice the true value is $\sigma(b)$ instead of $b$ for the first $k$ users. By Lemma 6.3, $\sigma(b_{k+1})$ has 0 confirmation probability. By Lemma 6.6, the total user utility under $\mathbf{s}$ is the same as under $\mathbf{s}^{\prime}=(\underbrace{\sigma(b),\ldots,\sigma(b)}_{k},\sigma(b-\delta))$ where $\delta>0$ can be arbitrarily small — since in both scenarios, the last bid is unconfirmed. Now, suppose the bid vector is $\mathbf{s}^{\prime}$. Fix any user with the true value $\sigma(b)$, say the first user. If it lowered its bid to anything less than $\sigma(b-\delta)$, its confirmation probability must be 00 by Lemma 6.3. Since the mechanism is UIC, by Myerson’s lemma, the payment of the first user is at least $\sigma(b-\delta)\cdot x_{1}(\mathbf{s}^{\prime})$. In this case, the first user’s utility is at most $(\sigma(b)-\sigma(b-\delta))\cdot x_{1}(\mathbf{s}^{\prime})\leq\sigma(b)-\sigma(b-\delta)$. Therefore, the total user utility under $\mathbf{s}^{\prime}$ (and hence under $\mathbf{s}$) is at most $k\cdot(\sigma(b)-\sigma(b-\delta))$. Since $\sigma$ is continuous at $b$, and the above holds for any $\delta>0$, the total user utility under $\mathbf{s}$ is at most 00. By the individual rationality of the payment rule, each user’s utility is non-negative when it bids truthfully. Therefore, the expected total user utility is 00.
@@
We next show that under $\mathbf{s}_{i-1}$, total user utility is 00. Consider another bid vector
$\mathbf{s}^{\prime}:=(\underbrace{\sigma(b),\ldots,\sigma(b)}_{i-1},\sigma(b-\delta),\ldots,\sigma(b-(k+2-i)\delta))$ for some $\delta>0$. By the argument above, the bids numbered $i+1,\ldots,k+1$ are not confirmed. By Lemma 6.6, the total user utility under $\mathbf{s}_{i-1}$ is the same as under the scenario $\mathbf{s}^{\prime}$. In scenario $\mathbf{s}^{\prime}$, if any of the $\sigma(b)$ bidders, say the first bidder, lowers its bid from $\sigma(b)$ to anything less than $\sigma(b-(k+2-i)\delta)$, its confirmation probability becomes 00 by Lemma 6.3. By UIC and Myerson, the first user’s payment under $\mathbf{s}^{\prime}$ is at least $\sigma(b-(k+2-i)\delta)\cdot x_{1}(\mathbf{s}^{\prime})$, and its utility is at most $(\sigma(b)-\sigma(b-(k+2-i)\delta))\cdot x_{1}(\mathbf{s}^{\prime})\leq\sigma(b)-\sigma(b-(k+2-i))\delta$. Therefore, the total user utility under $\mathbf{s}^{\prime}$ (and also under $\mathbf{s}_{i-1}$) is upper bounded by $k\cdot(\sigma(b)-\sigma(b-(k+2-i)\delta))$. Since $\sigma$ is continuous at $b$ and the above holds for any arbitrarily small $\delta>0$, it means that the total user utility under $\mathbf{s}_{i-1}$ is at most 00. By the individual rationality of the payment rule, each user’s utility is non-negative when it bids truthfully. Therefore, the expected total user utility is 00. ∎
@@
By Lemma 6.7, there exists a constant $c^{*}$ such that under the scenario $\mathbf{a}:=\left(\sigma(a^{*}),\sigma(a_{2}),\ldots,\sigma(a_{k+1})\right)$ where $a^{*}>a_{2}>\ldots>a_{k+1}>c^{*}$, the utility of the first user is 00 assuming it is bidding truthfully. Moreover, all users except the first user must have 0 confirmation probability. By Lemma 6.6, the total expected user utility under $\mathbf{a}$ and under a single bid $\sigma(a^{*})$ are the same. Therefore, under a single bid $\sigma(a^{*})$, the user’s utility is also 0. ∎
@@
We will show that under any sufficiently large $a$ at which $\sigma$ is continuous, the confirmation probability under a single bid $\sigma(a)$ is non-zero. If we can show this, then we can show a contradiction to UIC and the Myerson’s lemma. Specifically, consider $b>a$ which both are sufficiently large and at which $\sigma$ is continuous (Corollary 6.5 guarantees such $a,b$ exist). By Lemma 6.8, under both scenarios, the user utility is 00 (assuming the bid, $\sigma(a)$ or $\sigma(b)$, is the same as the user’s true value). By Lemma 6.4, for sufficiently large $a$, we have $\sigma(b)>\sigma(a)$. However, by Myerson’s lemma, if the confirmation probabilities under a single $\sigma(a)$ and under a single $\sigma(b)$ are both non-zero, it cannot be that in both cases the user’s utility is 00. Otherwise, the user with the true value $\sigma(b)$ can underbid $\sigma(a)$. Since the confirmation probability is non-zero and the payment is at most $\sigma(a)$, the user enjoys positive utility, which violates UIC.
@@
Remark 2.
Notice that Chung and Shi [CS23] showed that no TFM can simultaneously satisfy UIC and $1$-SCP when the block size is finite. Their impossibility does not rely on MIC, while MIC is necessary for proving Theorem 6.9. Specifically, consider the second-price auction without burning defined as follows. Given the block size $k$, the miner includes the top up to $k$ bids in the block. If the number of included bids is less than $k$, then all included bids are confirmed, and pay nothing. If the number of included bids is $k$, then the top $k-1$ bids are confirmed, while the $k$-th bid is unconfirmed (break ties arbitrarily). All confirmed bids pay the $k$-th bid, and all the payments go to the miner. This mechanism satisfies UIC since the confirmation and the payment satisfy the requirements of Myerson’s lemma. The mechanism also satisfies OCA-proofness when $\sigma$ is just bidding truthfully since the confirmed bids are always the ones with the highest true values, which maximizes the social welfare. However, it does not satisfy MIC, since the miner is incentivized to inject a fake bid to increase the $k$-th bid.
@@ 7 How to Circumvent the Impossibilities
In this section, we discuss possible ways to circumvent the impossibilities. Recall that in the definition of OCA-proofness, we require all users act independently in the globally optimal strategy $\sigma$, and $\sigma$ should only output a single bid. If we remove either of the two requirements, we can have mechanisms that satisfies UIC, MIC and OCA-proofness, where we present in Sections 7.1 and 7.2. These mechanisms are somewhat contrived and not obviously relevant for real-world blockchain protocols; the primary point of these mechanisms is to “stress-test” the definitions, and these examples highlight the subtlety of the modeling and help us to have a better understanding of the trade-offs between various notions of collusion-resilience. In Section 7.3, we consider an even weaker definition, called inclusion-rule-respecting, which might be the weakest but still meaningful definition that captures the notion of “no way to steal from the protocol.” Finally, in Section 7.4, we discuss the use of cryptography in the TFM literature, and point out a future direction by using cryptography and Bayesian incentive compatibility.

### v3

@@ 5.2 Formal Proofs
We now show that the confirmation probability under a single bid $a$ for a sufficiently large $a$ is non-zero assuming the miner honestly follows the inclusion rule. Suppose this is not true for the sake of a contradiction; that is, when the bid vector is just $a$ and the miner is honest, no one has positive confirmation probability, so the social welfare is zero. Since the mechanism is non-trivial, there must exist a scenario ${\mathbf{b}}=(b_{1},\dots,b_{t})$ such that some user $i$ has positive confirmation probability $p({\mathbf{b}})$. Let $a>\frac{1}{p({\mathbf{b}})}\cdot\sum_{j=1}^{t}b_{j}$. Then, imagine that there is only one user with true value $a$. The global coalition can replace the user’s primary bid with $b_{i}$, inject the fake bids ${\mathbf{b}}_{-i}$, and pretend that the bid vector is $\mathbf{b}$}. This way, the user would get positive confirmation probability $p({\mathbf{b}})$. Since its payment is at most $\sum_{j=1}^{t}b_{j}$, the user’s expected utility (and thus the social welfare) is positive. This violates global SCP.
@@ 6.1 A Non-Truthful Mechanism with UIC + MIC + OCA-Proof
The rationale of the design is to signal to the mechanism when everyone is adopting the globally optimal strategy $\sigma$ (as opposed to the bidding rule used to establish UIC). When the mechanism detects that everyone is behaving according to $\sigma$, it adopts a different behavior to optimize social welfare. We use the range $[0,1)$ to encode the actual bid, and use the range $1,\infty)$ for signalling. While the resulting mechanism is somewhat contrived and not necessarily meaningful from a practical point of view, it clarifies which notions of collusion-resilience most accurately capture the intended modeling goals and illustrates some technical challenges involved in the proof in Section 6.2. Consider the following TFM:
@@ 6.2 Impossibility of UIC + MIC + OCA-Proof for Truthful Mechanisms
The structure of the proof in this section is similar to the proof in Section 5. In fact, Lemmas 6.3, 6.4, 6.7, 6.8, 6.9 and 6.10 are the analogs of Lemmas 5.1, 5.2, 5.3, 5.4, 5.5 and 5.6, respectively, except that we need to work on the images of $\sigma$ in order to apply OCA-proofness. Before diving into the proof, we make a few remarks.
@@
  • As we have seen a feasibility for non-truthful mechanisms in Section 6.1, we must require the bidding rule to be truth-telling to reach an impossibility. Indeed, the proofs of Lemmas 6.5, 6.8, 6.9 and 6.10 rely on Myerson’s lemma, and Myerson’s lemma requires the mechanism to be truthful. Notice that the mechanism in Section 6.1 does not satisfy the requirements of Myerson’s lemma. If we apply the revelation principle to make the mechanism become truthful so that Myerson’s lemma could apply, it breaks OCA-proofness (see Remark 6.2).
@@
Imagine that $k+1$ players each have true values $b_{1},\ldots,b_{k+1}$, and they all bid according to the globally optimal strategy $\sigma$. It suffices to show the following: suppose $\sigma(b_{k+1})$ has a non-zero probability $p$ of being included in the block, then conditioned on being included, its probability of confirmation must be 0. Suppose this is not true, and conditioned on being included, $\sigma(b_{k+1})$ has probability $p^{\prime}>0$ of being confirmed. Then, the global coalition can adopt the following OCA: run the honest inclusion rule in its head, including flipping any coins required by the inclusion rule, and let $\bf B$} denote the resulting block. If $\sigma(b_{k+1})$ is not included in $\bf B$}, simply output the block $\bf B$}. Else if $\sigma(b_{k+1})$ is included in $\bf B$}, then let $\sigma(b_{i})$ be some bid that is left out — such a bid must exist because the block size is only $k$. Now, have the $i$-th user bid $\sigma(b_{k+1})$ and make an actual block $\bf B$} where the bid $\sigma(b_{k+1})$ corresponds to user $i$’s bid. The expected increase in social welfare relative to everyone adopting $\sigma$ is $p\cdot p^{\prime}\cdot(b_{i}-b_{k+1})>0$, which violates OCA-proofness and the fact that $\sigma$ is the social-welfare-maximizing strategy. ∎
@@
A well-known fact is that any function that is monotone must have only countably many discontinuities within any interval $x_{1},x_{2}]$ where $x_{2}>x_{1}$; further, all such discontinuities must be jump discontinuities. By Lemma 6.5, there exists a constant $c^{*}$ such that $\sigma$ is monotone for all inputs larger than $c^{*}$. Thus, for any $a>c^{*}$, within the interval $[a,a+1]$, $\sigma$ is monotone and there must exist a point $a^{*}$ such that $\sigma$ is continuous at $a^{*}$. ∎
@@
  • Moreover, if there are $k+1$ users with the true value $\mathbf{s}$ and $\sigma$ is continuous at $b$, the expected total user utility is 0 when all users bid truthfully.
@@
The base case is when $i=k$; that is, there are $k+1$ users with the true values $\mathbf{s}=(\underbrace{\sigma(b),\ldots,\sigma(b)}_{k},\sigma(b_{k+1}))$. Notice the true value is $\sigma(b)$ instead of $b$ for the first $k$ users. By Lemma 6.4, $\sigma(b_{k+1})$ has 0 confirmation probability. By Lemma 6.7, the total user utility under $\mathbf{s}$ is the same as under $\mathbf{s}^{\prime}=(\underbrace{\sigma(b),\ldots,\sigma(b)}_{k},\sigma(b-\delta))$ where $\delta>0$ can be arbitrarily small — since in both scenarios, the last bid is unconfirmed. Now, suppose the bid vector is $\mathbf{s}^{\prime}$. Fix any user with the true value $\sigma(b)$, say the first user. If it lowered its bid to anything less than $\sigma(b-\delta)$, its confirmation probability must be 0 by Lemma 6.4. Since the mechanism is UIC, by Myerson’s lemma, the payment of the first user is at least $\sigma(b-\delta)\cdot x_{1}(\mathbf{s}^{\prime})$. In this case, the first user’s utility is at most $(\sigma(b)-\sigma(b-\delta))\cdot x_{1}(\mathbf{s}^{\prime})\leq\sigma(b)-\sigma(b-\delta)$. Therefore, the total user utility under $\mathbf{s}^{\prime}$ (and hence under $\mathbf{s}$) is at most $k\cdot(\sigma(b)-\sigma(b-\delta))$. Since $\sigma$ is continuous at $b$, and the above holds for any $\delta>0$, the total user utility under $\mathbf{s}$ is at most 0. By the individual rationality of the payment rule, each user’s utility is non-negative when it bids truthfully. Therefore, the expected total user utility is 0.
@@
We next show that under $\mathbf{s}_{i-1}$, total user utility is 0. Consider another bid vector
$\mathbf{s}^{\prime}:=(\underbrace{\sigma(b),\ldots,\sigma(b)}_{i-1},\sigma(b-\delta),\ldots,\sigma(b-(k+2-i)\delta))$ for some $\delta>0$. By the argument above, the bids numbered $i+1,\ldots,k+1$ are not confirmed. By Lemma 6.7, the total user utility under $\mathbf{s}_{i-1}$ is the same as under the scenario $\mathbf{s}^{\prime}$. In scenario $\mathbf{s}^{\prime}$, if any of the $\sigma(b)$ bidders, say the first bidder, lowers its bid from $\sigma(b)$ to anything less than $\sigma(b-(k+2-i)\delta)$, its confirmation probability becomes 0 by Lemma 6.4. By UIC and Myerson, the first user’s payment under $\mathbf{s}^{\prime}$ is at least $\sigma(b-(k+2-i)\delta)\cdot x_{1}(\mathbf{s}^{\prime})$, and its utility is at most $(\sigma(b)-\sigma(b-(k+2-i)\delta))\cdot x_{1}(\mathbf{s}^{\prime})\leq\sigma(b)-\sigma(b-(k+2-i))\delta$. Therefore, the total user utility under $\mathbf{s}^{\prime}$ (and also under $\mathbf{s}_{i-1}$) is upper bounded by $k\cdot(\sigma(b)-\sigma(b-(k+2-i)\delta))$. Since $\sigma$ is continuous at $b$ and the above holds for any arbitrarily small $\delta>0$, it means that the total user utility under $\mathbf{s}_{i-1}$ is at most 0. By the individual rationality of the payment rule, each user’s utility is non-negative when it bids truthfully. Therefore, the expected total user utility is 0. ∎
@@
By Lemma 6.8, there exists a constant $c^{*}$ such that under the scenario $\mathbf{a}:=\left(\sigma(a^{*}),\sigma(a_{2}),\ldots,\sigma(a_{k+1})\right)$ where $a^{*}>a_{2}>\ldots>a_{k+1}>c^{*}$, the utility of the first user is 0 assuming it is bidding truthfully. Moreover, all users except the first user must have 0 confirmation probability. By Lemma 6.7, the total expected user utility under $\mathbf{a}$ and under a single bid $\sigma(a^{*})$ are the same. Therefore, under a single bid $\sigma(a^{*})$, the user’s utility is also 0. ∎
@@
We will show that under any sufficiently large $a$ at which $\sigma$ is continuous, the confirmation probability under a single bid $\sigma(a)$ is non-zero. If we can show this, then we can show a contradiction to UIC and the Myerson’s lemma. Specifically, consider $b>a$ which both are sufficiently large and at which $\sigma$ is continuous (Corollary 6.6 guarantees such $a,b$ exist). By Lemma 6.9, under both scenarios, the user utility is 0 (assuming the bid, $\sigma(a)$ or $\sigma(b)$, is the same as the user’s true value). By Lemma 6.5, for sufficiently large $a$, we have $\sigma(b)>\sigma(a)$. However, by Myerson’s lemma, if the confirmation probabilities under a single $\sigma(a)$ and under a single $\sigma(b)$ are both non-zero, it cannot be that in both cases the user’s utility is 0. Otherwise, the user with the true value $\sigma(b)$ can underbid $\sigma(a)$. Since the confirmation probability is non-zero and the payment is at most $\sigma(a)$, the user enjoys positive utility, which violates UIC.
@@
Remark 6.11.
Notice that Chung and Shi [CS23] showed that no TFM can simultaneously satisfy UIC and $1$-SCP when the block size is finite. Their impossibility does not rely on MIC, while MIC is necessary for proving Theorem 6.10. Specifically, consider the second-price auction without burning defined as follows. Given the block size $k$, the miner includes the top up to $k$ bids in the block. If the number of included bids is less than $k$, then all included bids are confirmed, and pay nothing. If the number of included bids is $k$, then the top k1k-1 bids are confirmed, while the $k$-th bid is unconfirmed (break ties arbitrarily). All confirmed bids pay the $k$-th bid, and all the payments go to the miner. This mechanism satisfies UIC since the confirmation and the payment satisfy the requirements of Myerson’s lemma. The mechanism also satisfies OCA-proofness when $\sigma$ is just bidding truthfully since the confirmed bids are always the ones with the highest true values, which maximizes the social welfare. However, it does not satisfy MIC, since the miner is incentivized to inject a fake bid to increase the $k$-th bid.
@@ 7 How to Circumvent the Impossibilities
In this section, we discuss possible ways to circumvent the impossibilities. Recall that in the definition of OCA-proofness, we require all users act independently in the globally optimal strategy $\sigma$, and $\sigma$ should only output a single bid. If we remove either of the two requirements, we can have mechanisms that satisfies UIC, MIC and OCA-proofness, where we present in Sections 7.1 and 7.2. These mechanisms are somewhat contrived and not obviously relevant for real-world blockchain protocols; the primary point of these mechanisms is to “stress-test” the definitions, and these examples highlight the subtlety of the modeling and help us to have a better understanding of the trade-offs between various notions of collusion-resilience. In Section 7.3, we consider an even weaker definition, called inclusion-rule-respecting, which might be the weakest but still meaningful definition that captures the notion of “no way to steal from the protocol.” Later in Section 8, we introduce a new model that captures the use of trusted hardware and formalize a Bayesian notion of incentive compatibility. We show that, in this model, it is possible to construct a TFM that satisfies Bayesian incentive compatibility.